# Hanging on the cliff: EMRI formation with local two-body relaxation and post-Newtonian dynamics

### v1

@@ 4.4 Effect on EMRI rates
Despite cliffhangers becoming increasingly more common at lower MBH masses, their effect on the overall EMRI rates gets progressively less important. This is mainly because the peak energy of the loss cone flux, $\mathcal{E}_{\rm lc,p}$, becomes comparable (and even bigger) then $\mathcal{E}_{\mathrm{c}}$ at low MBH masses. In the $M_{\bullet}=10^{5}\,{\rm M}_{\odot}$ case, $\mathcal{E}_{\mathrm{c}}\gtrsim\mathcal{E}_{\rm lc,p}$, therefore adding a significant population of cliffhanger EMRIs at lower energies has a relevant impact on the overall EMRI population. In fact, despite the overall EMRI rate being substantially unaffected (Table 1) we find that ${\sim}37\%$ of those EMRIs are cliffhanger and that the peak of the EMRI flux occurs at $a>a_{\mathrm{c}}$ ($\mathcal{E}<\mathcal{E}_{\mathrm{c}}$) which might result in extremely eccentric systems. Conversely, the overall EMRI rate decreases in the $M_{\bullet}=10^{4}\,{\rm M}_{\odot}$ case, with cliffhangers accounting only to ${\sim}10\%$ of the overall EMRI population. Notably, for such a low MBH mass, $\mathcal{E}_{\mathrm{c}}<\mathcal{E}_{\rm lc,p}$ meaning that most EMRIs form well inside $a_{\mathrm{c}}$ potentially resulting in lower eccentric systems.
This analysis shows that while the impact of the detailed treatment of nuclear dynamics and the appearance of cliffhangers on the EMRI rate might be only moderate, the energy distributon of EMRIs (and hence their detailed properties) can be highly affected. In particular, a simplistic treatment of EMRIs as systems mostly forming at $a=a_{\mathrm{c}}$ fails in capturing the rich and MBH mass dependent features of this class of GW sources.
@@ 5 Discussion and conclusions
In this work, we developed a formalism to locally account for the effects of two-body relaxation on the formation of EMRIs and DPs around an MBH at the core of a nuclear cluster, for which we considered a two-population model. We examined a binary system comprised of an MBH and a stellar-mass BH orbiting around it, evolving such system by using PN dynamics up to the 2.5PN term. We ran a number of simulations and studied the EMRI-to-DP ratio as a function of the initial semi-major axis $a_{0}$ of the orbit through the examination of the function $S(a_{0})$, for five different central MBH masses. Our main findings can be summarised as follows.
@@
    For low MBH masses, the classical picture for which there is a critical value $a_{\mathrm{c}}$ such that EMRIs only form at $a_{0}\ll a_{\mathrm{c}}$, while no EMRI occur at $a_{0}\gg a_{\mathrm{c}}$, breaks down. Instead, we find a significant population of EMRIs forming for large values of $a_{0}$. Such population is consistent with the novel phenomenon of cliffhanger EMRIs, first identified in QS23.
@@
    Cliffhanger EMRIs are more common for low-mass MBHs because there is a larger region in the $(1-e,a)$ phase space which allows for a drastic shrinking of the orbit in a single pericentre passage due to GW radiation. Following this analytical argument, we predict the maximum MBH mass expected to produce cliffhanger EMRIs given our assumption to be roughly $3\times 10^{5}\,\mathrm{M_{\sun}}$, which was verified in our numerical investigation.
@@
    The choice of locally accounting for two-body relaxation does significantly influence the EMRI-to-DP ratio for $M_{\bullet}\lesssim 10^{5}\,\mathrm{M_{\sun}}$, leading to a much larger fraction of cliffhanger EMRIs compared to what found in the orbit-averaged approximation.
@@
Such results might have a significant impact on the estimates of EMRI detection rates by LISA. Currently, these estimates are still highly uncertain, ranging from a few to a few thousands of EMRIs detected per year (Babak et al. 2017), and do not account for the novel population of cliffhanger EMRIs. Furthermore, we note that many cliffhanger EMRIs get to the GW emission region while grazing the loss cone edge, which might imply a significant sub-population of extremely eccentric EMRIs in the LISA band. We defer the investigation of observable cliffhanger EMRI properties by LISA to future work.
@@
Interestingly, we observed that our results are quite sensitive to the choice of the loss cone radius $R_{\mathrm{lc}}$. Indeed, as explained in Sec. 4.3 part of the reason for the differences between the $S(a_{0})$ obtained with Newtonian and PN dynamics likely stems from switching from $R_{\mathrm{lc}}=8R_{\mathrm{g}}$ to $R_{\mathrm{lc}}=5.6R_{\mathrm{g}}$, as plunges are more difficult to form and the cliffhanger region grows. For this reason, we believe that the closest approach to a BH on an eccentric orbit should be better explored at PN order, similarly to how it has been already done for circular orbits (e.g. Schäfer & Jaranowski 2024). Compared to Schwarzschild MBHs, spinning MBHs present innermost stable orbits which are closer or further away, respectively if the orbit is prograde or retrograde. This again influences the loss cone radius, and consequently the EMRI-to-DP ratio. For this reason, we plan on updating our code to introduce spin effects.
@@
Finally, despite improving on realism by not employing orbit-averages in accounting for two-body relaxation, our approach is still limited on some fronts. Our analysis currently addresses spherical systems with a central Schwarzshild MBH. However, these results might not hold for axisymmetric or triaxial systems due to aspherical two-body relaxation and orbits naturally penetrating the loss cone (Vasiliev & Merritt 2013; Kaur & Stone 2024), or spinning MBHs (Amaro-Seoane et al. 2013). Moreover, even restricting to the class of systems we are investigating, we explore a limited region of the $(1-e,a)$ plane. Since this choice is not biased (at least for low MBH masses) towards EMRIs or DPs, this does not influence our results on the EMRI-to-DP ratio, but it requires the introduction of corrections on the counts of EMRIs for large MBH masses and limits the maximum $a_{0}$ which we can explore. The latter statement is particularly true for low MBH masses, for which we are not able to reach $a_{0}\approx R_{\mathrm{inf}}$, as most simulated orbits starting from large $a_{0}$ result in reinitialisations, thus requiring a very large number of runs to find the EMRI-to-DP ratio within a reasonably small error. We plan on solving this issue in a follow-up work, as pushing the exploration to larger $a_{0}$ values is fundamental to understand whether cliffhanger EMRIs make $S(a_{0})$ reach a plateau $S(\infty)\neq 0$ or not.
@@ Appendix A Diffusion coefficients
Here $\ln{\Lambda}$ is the Coulomb logarithm, which we set to $10$ as an order of magnitude approximation (Spinnato et al. 2003; Merritt 2013b), while the functions $\mathcal{I}_{k,\alpha}$ and $\mathcal{J}_{k,\alpha}$ are defined as:
@@ Appendix B From $\delta v_{\parallel}$ and $\delta v_{\perp}$ to $\delta\mathcal{E}$ and $\delta J$
from which follows that $J=|\@vec{J}|$ changes as:
@@
$=J\left(\frac{\delta v_{\parallel}}{v}+\delta v_{\perp}\frac{v_{\mathrm{r}}}{vv_{\mathrm{t}}}\sin{\theta}+\delta v_{\parallel}\delta v_{\perp}\frac{v_{\mathrm{r}}}{vv_{\mathrm{t}}}\sin{\theta}+\frac{1}{2}\left(\frac{\delta v_{\perp}\cos{\theta}}{v_{\mathrm{t}}}\right)^{2}\right)$
$+\mathcal{O}\left((\delta v)^{3}\right)\,,$
@@
More detailed steps can be found in Broggi (2024).
@@ Appendix C From $\Delta\mathcal{E}$ and $\Delta J$ to $\Delta v$
where $\beta$ is the angle between vectors ${J_{1}}$ and ${J}$. Using the relations found above, we can rewrite the system as:
@@
Finally, after a sequence of substitutions, the system above can be written as:

### v2

@@ 4.4.2 Effects on EMRIs from lower central MBH masses
Despite cliffhangers becoming increasingly more common at lower MBH masses, their effect on the overall EMRI rates gets progressively less important. This is mainly because the peak energy of the loss cone flux, $\mathcal{E}_{\rm lc,p}$, becomes comparable (and even bigger) then $\mathcal{E}_{\mathrm{c}}$ at low MBH masses. In the $M_{\bullet}=10^{5}\,{\rm M}_{\odot}$ case, $\mathcal{E}_{\mathrm{c}}\gtrsim\mathcal{E}_{\rm lc,p}$, therefore adding a significant population of cliffhanger EMRIs at lower energies has a relevant impact on the overall EMRI population. Despite the overall EMRI rate being substantially unaffected (Table 1) we find that ${\sim}37\%$ of those EMRIs are cliffhanger and that the peak of the EMRI flux occurs at $a>a_{\mathrm{c}}$ ($\mathcal{E}<\mathcal{E}_{\mathrm{c}}$) which might result in extremely eccentric systems. Conversely, the overall EMRI rate decreases in the $M_{\bullet}=10^{4}\,{\rm M}_{\odot}$ case, with cliffhangers accounting only to ${\sim}10\%$ of the overall EMRI population. Notably, for such a low MBH mass, $\mathcal{E}_{\mathrm{c}}<\mathcal{E}_{\rm lc,p}$ meaning that most EMRIs form well inside $a_{\mathrm{c}}$ potentially resulting in lower eccentric systems.
This analysis shows that while the impact of the detailed treatment of nuclear dynamics and the appearance of cliffhangers on the EMRI rate might be only moderate, the energy distribution of EMRIs (and hence their detailed properties) can be highly affected. In particular, a simplistic treatment of EMRIs as systems mostly forming at $a=a_{\mathrm{c}}$ fails in capturing the rich and MBH mass dependent features of this class of GW sources.
@@ 5 Discussion and conclusions
In this work we developed a formalism to locally account for the effects of two-body relaxation on the formation of EMRIs and DPs around an MBH at the core of a nuclear cluster, for which we considered a two-population model. We examined a binary system comprised of an MBH and a stellar-mass BH orbiting around it, evolving the system by using PN dynamics up to the 2.5PN term. We ran a number of simulations and studied the EMRI-to-DP ratio as a function of the initial semi-major axis $a_{0}$ of the orbit through the examination of the function $S(a_{0})$, for five different central MBH masses. Our main findings can be summarised as follows:
@@
    For low MBH masses the classical picture, for which there is a critical value $a_{\mathrm{c}}$ such that EMRIs only form at $a_{0}\ll a_{\mathrm{c}}$ while no EMRIs occur at $a_{0}\gg a_{\mathrm{c}}$, breaks down. Instead, we find a significant population of EMRIs forming for large values of $a_{0}$. Such a population is consistent with the novel phenomenon of cliffhanger EMRIs, first identified in QS24.
@@
    Cliffhanger EMRIs are more common for low-mass MBHs because there is a larger region in the $(1-e,a)$ phase space which allows for a drastic reduction of the orbit in a single pericentre passage due to GW radiation. Following this analytical argument, we predict that the maximum MBH mass expected to produce cliffhanger EMRIs, given our assumption, is roughly $3\times 10^{5}\,\mathrm{M_{\sun}}$, which was verified in our numerical investigation.
@@
    The choice of locally accounting for two-body relaxation does significantly influence the EMRI-to-DP ratio for $M_{\bullet}\lesssim 10^{5}\,\mathrm{M_{\sun}}$, leading to a much larger fraction of cliffhanger EMRIs compared to that found in the orbit-averaged approximation.
@@
These results might have a significant impact on the estimates of EMRI detection rates by LISA. Currently, these estimates are still highly uncertain, ranging from a few to a few thousand EMRIs detected per year (Babak et al. 2017), and do not account for the novel population of cliffhanger EMRIs. Furthermore, we note that many cliffhanger EMRIs get to the GW emission region while grazing the loss cone edge, which might imply a significant sub-population of extremely eccentric EMRIs in the LISA band. We defer the investigation of observable cliffhanger EMRI properties by LISA to future work.
@@
Interestingly, we observed that our results are quite sensitive to the choice of the loss cone radius $R_{\mathrm{lc}}$. As explained in Sect. 4.3, part of the reason for the differences between the $S(a_{0})$ obtained with Newtonian and PN dynamics likely stems from switching from $R_{\mathrm{lc}}=8R_{\mathrm{g}}$ to $R_{\mathrm{lc}}=5.6R_{\mathrm{g}}$, as plunges are more difficult to form and the cliffhanger region grows. For this reason, we believe that the closest approach to a BH on an eccentric orbit should be better explored at PN order, similarly to how it has already been done for circular orbits (e.g. Schäfer & Jaranowski 2024). Compared to Schwarzschild MBHs, spinning MBHs present innermost stable orbits which are closer or further away, respectively, depending on whether the orbit is prograde or retrograde. This again influences the loss cone radius, and consequently the EMRI-to-DP ratio. For this reason, we plan on updating our code to introduce spin effects.
Despite improving on realism by not employing orbit-averages in accounting for two-body relaxation, our approach is still limited on some fronts. Our analysis currently addresses spherical systems with a central Schwarzshild MBH. However, these results might not hold for axisymmetric or triaxial systems due to aspherical two-body relaxation and orbits naturally penetrating the loss cone (Vasiliev & Merritt 2013; Kaur & Stone 2024), or for spinning MBHs (Amaro-Seoane et al. 2013). Furthermore, the two-component Bahcall-Wolf profile we assumed might underestimate the number of stellar-mass BHs for $r<R_{\mathrm{inf}}$, as previous works (Hopman & Alexander 2006; Broggi et al. 2022; Rom et al. 2024) have found more segregated profiles. As a result, in our simulations stars dominate two-body relaxation over the entire $r$ range, while for a completely relaxed system stellar-mass BHs should dominate for $r\lesssim 0.1R_{\mathrm{inf}}$. While this could affect the shape of $S(a_{0})$, and in particular the value of $a_{\mathrm{c}}$, we note that for the EMRI rate computations of Sect. 4.4 the density profiles were consistently relaxed using the code presented in Broggi et al. (2022), thus resulting in a more segregated distribution of stellar-mass BHs.
@@
Finally, even restricting the discussion to the class of systems we are investigating, we explored a limited region of the $(1-e,a)$ plane. Since this choice was not biased (at least for low MBH masses) towards EMRIs or DPs, this did not influence our results on the EMRI-to-DP ratio, but it required the introduction of corrections on the counts of EMRIs for high MBH masses and limits the maximum $a_{0}$ that we can explore. This is particularly true for low MBH masses, for which we are not able to reach $a_{0}\approx R_{\mathrm{inf}}$, as most simulated orbits starting from large $a_{0}$ result in reinitialisations, thus requiring a very large number of runs to find the EMRI-to-DP ratio within a reasonably small error. We plan on solving this issue in a follow-up work, as pushing the exploration to larger $a_{0}$ values is fundamental in order to understand whether cliffhanger EMRIs allow $S(a_{0})$ to reach a plateau $S(\infty)\neq 0$ or not.
@@ Appendix A Diffusion coefficients
Here $\ln{\Lambda}$ is the Coulomb logarithm, which we set to $10$ as an order of magnitude approximation (Spinnato et al. 2003; Merritt 2013b), while the functions $\mathcal{I}_{k,\alpha}$ and $\mathcal{J}_{k,\alpha}$ are defined as
@@ Appendix B From $\delta v_{\parallel}$ and $\delta v_{\perp}$ to $\delta\mathcal{E}$ and $\delta J$
from which follows that $J=|\@vec{J}|$ changes as
@@
$=J\left(\frac{\delta v_{\parallel}}{v}+\delta v_{\perp}\frac{v_{\mathrm{r}}}{vv_{\mathrm{t}}}\sin{\theta}+\frac{1}{2}\left(\frac{\delta v_{\perp}\cos{\theta}}{v_{\mathrm{t}}}\right)^{2}\right)+\mathcal{O}\left((\delta v)^{3}\right)\,,$
@@
More detailed steps can be found in Broggi (2024). [^14] The expression for $\delta J$ reported in Broggi (2024) mistakenly reports an extra term proportional to $\delta v_{\parallel}\delta v_{\perp}$.
@@ Appendix C From $\Delta\mathcal{E}$ and $\Delta J$ to $\Delta v$
where $\beta$ is the angle between vectors ${J_{1}}$ and ${J}$. Using the relations found above, we can rewrite the system as
@@
Finally, after a sequence of substitutions, the system above can be written as